Light of wavelength 590 nm illuminates a diffraction grating. The second-order maximum is at angle 40.5

How many lines per millimeter does this grating have?
Answer #1

here,

let the line seperation is d

Now, for the second maximum

d * sin(theta) = m * wavelength

d * sin(40.5 degree) = 2 * 590 *10^-9

d = 1.816 *10^-6 m = 1.816 *10^-3 mm

lines per mm = 1/d

lines per mm = 1/(1.816 *10^-3)

lines per mm = 551

the lines per mm is 551 lines/mm